OpenStudy (anonymous):

Compute the slope of the secant line between the given x-values. f(x) = x^3 - x. a. x = 1 and x = 2 b. x = 2 and x = 3 c. x = 1.5 and x = 2 d. x = 2 and x = 2.5 e. x = 1.9 and x = 2 f. x = 2 and x = 2.1 g. Use parts a-f to estimate the slope of the tangent line at x = 2

OpenStudy (anonymous):

@satellite73 can u help?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[f(x)=x^3-x\] so \[f(2)=2^3-2=6\] the point \((2,6)\) is on the graph \[f(1)=1^3-1=0\] the point \((1,0)\) is on the graph

OpenStudy (anonymous):

the slope of the line between those two points is \[\frac{6-0}{2-1}=6\]

OpenStudy (anonymous):

okay got you

OpenStudy (anonymous):

now you have a lot of questions to do, and i have to run, but the idea is the same for all of them

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

so for example x = 1.9 and x = 2 you want to compute \[\frac{f(2)-f(1.9)}{2-1.9}\]

OpenStudy (anonymous):

i would use a calculator to compute \(f(1.9)=(1.9)^2-1.9\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

1.71

OpenStudy (anonymous):

thanks i'll do the rest

OpenStudy (anonymous):

you get \[\frac{6-4.959}{.1}=\frac{1.041}{.1}=10.41\] i think check good luck

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hey umm so how about g? what does it say to do?